Show that each of these conditional statements is a tautology by using truth tables.

 

a) (p q) → p                b) p → (p q)

c) p → (p q)                            d) (p q) → (p q)

e) ¬(p q) → p                        f) ¬(p q) → ¬q



1
Expert's answer
2021-12-06T16:24:00-0500

Let us show that each of these conditional statements is a tautology by using truth tables.

 

a) (∧ q) → p               


"\begin{array}{||c|c||c|c||} 
\hline\hline 
p & q &p ∧ q & (p ∧ q) → p \\ 
\hline\hline 
0 & 0 & 0 & 1\\ 
\hline 
0 & 1 & 0 & 1
 \\ \hline 
1 & 0 & 0 & 1\\ 
\hline 
1 & 1 & 1 & 1\\ 
\hline\hline 
\end{array}"


Since the last column contains only 1, we conclude that this formula is a tautology.


b) → (∨ q)


"\begin{array}{||c|c||c|c||} 
\hline\hline 
p & q &p \lor q &p → (p ∨ q) \\ 
\hline\hline 
0 & 0 & 0 & 1\\ 
\hline 
0 & 1 & 1 & 1
 \\ \hline 
1 & 0 & 1 & 1\\ 
\hline 
1 & 1 & 1 & 1\\ 
\hline\hline 
\end{array}"


Since the last column contains only 1, we conclude that this formula is a tautology.


c) → (→ q)                            


"\begin{array}{||c|c||c|c|c||} 
\hline\hline 
p & q &\neg p & p \to q &\neg p → (p \to q) \\ 
\hline\hline 
0 & 0 & 1 & 1 & 1\\ 
\hline 
0 & 1 & 1 & 1 & 1
 \\ \hline 
1 & 0 & 0 & 0 & 1\\ 
\hline 
1 & 1 & 0 & 1 &1\\ 
\hline\hline 
\end{array}"


Since the last column contains only 1, we conclude that this formula is a tautology.


d) (∧ q) → (→ q)


"\begin{array}{||c|c||c|c|c||} 
\hline\hline 
p & q &p ∧ q & p \to q &(p ∧ q) → (p → q)\\ 
\hline\hline 
0 & 0 & 0 & 1 & 1\\ 
\hline 
0 & 1 & 0 & 1 & 1
 \\ \hline 
1 & 0 & 0 & 0 & 1\\ 
\hline 
1 & 1 & 1 & 1 &1\\ 
\hline\hline 
\end{array}"


Since the last column contains only 1, we conclude that this formula is a tautology.


e) ¬(→ q) → p                        


"\begin{array}{||c|c||c|c|c||} 
\hline\hline 
p & q & p \to q& \neg (p → q) & \neg(p → q) → p \\ 
\hline\hline 
0 & 0 & 1 & 0 & 1\\ 
\hline 
0 & 1 & 1 & 0 & 1
 \\ \hline 
1 & 0 & 0 & 1 & 1\\ 
\hline 
1 & 1 & 1 & 0 &1\\ 
\hline\hline 
\end{array}"


Since the last column contains only 1, we conclude that this formula is a tautology.


f) ¬(→ q) → ¬q


"\begin{array}{||c|c||c|c|c|c||} 
\hline\hline 
p & q & p \to q& \neg (p → q) & \neg q & \neg (p → q) \to \neg q\\ 
\hline\hline 
0 & 0 & 1 & 0 & 1 & 1\\ 
\hline 
0 & 1 & 1 & 0 & 0 & 1
 \\ \hline 
1 & 0 & 0 & 1 & 1 & 1\\ 
\hline 
1 & 1 & 1 & 0 & 0 & 1\\ 
\hline\hline 
\end{array}"


Since the last column contains only 1, we conclude that this formula is a tautology.